# Accurate Estimation of Mutual Information in High Dimensional Data

### v1

@@ A.7 Additional Figures
Figure 10: The KSG estimator does not properly estimate MI for high-dimensional ($K=500$) data even with a low-dimensional latent ($K_{Z}=10$) space. The KSG estimator [21] with code adapted from [31], was used to estimate information contained in $K=500$-dimensional data generated by teacher networks with latent dimensionality $K_{Z}=10$ and true MI of 4 bits, as used in Fig. 6. Left and right panels correspond to the undersampled ($N=256$) and the well-sampled ($N=2^{14}$) regimes. Best practices, which are similar to those used in Fig. 6, were used to estimate the MI with KSG estimator [40] at different numbers of nearest neighbors $k$. Specifically, as for NN estimators, data was partitioned into $\gamma$ non-overlapping partitions and estimates for each partition were found. The mean information and standard deviation using $\gamma$ partitions are plotted versus the number of partitions at different values of $k$. The KSG estimator does not asymptotically approach the correct value of information at $\gamma=0$, corresponding to the infinite data limit, and linear extrapolations (done the same way as in Fig. 6) is unreliable. Best practice is to not use the estimator in this case [40]. Indeed, in the undersampled regime, the estimator would predict MI of $1.8$, $1.3$, and $1.0$ bits for $k=1$, $5$, and $10$ respectively. For the better-sampled regime, the values would be $2.8$, $2.4$, $2.3$, $2.2$, and $2.1$ bits for $k=1$, $5$, $10$, $15$ and $20$ respectively. All these values are well below the ground truth of 4 bits: KSG is, indeed, failing. In contrast, neural estimators produce near-perfect results in both regimes (see Fig. 6).
Figure 11: Sample pairs from the Noisy MNIST dataset. Each pair $(X,Y)$ shares the same digit class label but consists of distinct, non-overlapping digit instances. The $X$ image is generated by applying a random rotation (uniform between 00 and $\pi/2$) and a random scaling (uniform between $0.5$ and $1.5$) to an MNIST digit. The $Y$ image is formed by applying a new instant of Perlin noise background to each digit, with the noise weight uniformly drawn from $[0,1]$. Both images are normalized to the $[0,1)$ intensity range and flattened to $784$-dimensional vectors. We generate up to $2^{18}$ unique training pairs from the MNIST training set and $1024$ from the test set, ensuring an approximately uniform digit distribution. These data preserve only the class-level semantic information ($I(X;Y)\approx\log_{2}10$ bits), providing a high-dimensional test case for mutual information estimation.

### v2

@@ A.8 Additional Figures
Figure 11: Sample pairs from the Noisy MNIST dataset. Each pair $(X,Y)$ shares the same digit class label but consists of distinct, non-overlapping digit instances. The $X$ image is generated by applying a random rotation (uniform between 0 and $\pi/2$) and a random scaling (uniform between $0.5$ and $1.5$) to an MNIST digit. The $Y$ image is formed by applying a new instant of Perlin noise background to each digit, with the noise weight uniformly drawn from $[0,1]$. Both images are normalized to the $[0,1)$ intensity range and flattened to $784$-dimensional vectors. We generate up to $2^{18}$ unique training pairs from the MNIST training set and $1024$ from the test set, ensuring an approximately uniform digit distribution. These data preserve only the class-level semantic information ($I(X;Y)\approx\log_{2}10$ bits), providing a high-dimensional test case for mutual information estimation.